# Enhancing Corpus Callosum Segmentation in Fetal MRI via Pathology-Informed Domain Randomization

### v1

@@ Abstract
Keywords:
Corpus callosum Segmentation Domain Randomization Deep Learning Data Augmentation.
@@ 2.1 Datasets
This study includes a total of 370 fetal subjects sourced from both private clinical cohorts and publicly available datasets. An overview of the datasets is provided in Table 1; we refer readers to the original publications for full details on acquisition protocols and annotation procedures. All data were anonymized and approved by the appropriate ethics committees prior to inclusion.111 KISPI: Ethical Committee of the Canton of Zurich, Switzerland (Decision numbers: 2017-00885, 2016-01019, 2017-00167). CHUV: Approved by the ethics committee of the Canton de Vaud (decision number CERVD 2021 00124). dHCP: Approved by the UK Health Research Authority (REC: 14/LO/1169), with written parental consent. For each subject, we used a super-resolution reconstructed T2-weighted MRI, resampled to an isotropic resolution of 0.5×0.5×0.5mm3 and padded or cropped to a fixed size of 256×256×256 voxels. For the KISPI, dHCP, and STA datasets, annotations were harmonized into a consistent 8-class labeling scheme: cerebrospinal fluid (CSF), grey matter (GM), white matter (WM), ventricles (VM), cerebellum (CBM), deep grey matter (SGM), brainstem (BSM), and CC. Details of the harmonization procedure are provided in the Supplementary Section S1.
@@
dHCP [28] Philips Achieva (3T) MB-TSE TR/TE: 2265/250 ms 1.1×1.1×2.2 mm³ IRTK [26]; [29] $0.8^{3}$ 21–38 248/0 dHCP [30]
STA [31] Siemens Skyra / Trio (3T) Philips Achieva (1.5T) SS-FSE TR/TE: 1400–2000/100–120 ms 0.9–1.1×0.9–1.1×2.0 mm³ SR1 [32] $1^{3}$ 21–38 18/0 FeTA [24]+ CC [27]
@@ 2.2 Pathology-Informed Domain Randomization for CCD
To ensure plausibility, an expert neuroradiologist reviewed the generated segmentations, both individual alterations and their combinations, to confirm broad anatomical constraints (for example, maintaining realistic ventricular dilation limits and proper continuity in simulations of pontocerebellar hypoplasia). However, the expert did not assess or correct finer-grained details (such as subtle changes in adjacent structures like the cingulate gyrus), both because detailed segmentations of these regions were not available and because enforcing such precision would conflict with the goal of domain randomization.

### v2

@@ 2.1 Datasets
This study includes a total of 370 fetal subjects sourced from both private clinical cohorts and publicly available datasets. An overview of the datasets is provided in Table 1; we refer readers to the original publications for full details on acquisition protocols and annotation procedures. All data were anonymized and approved by the appropriate ethics committees prior to inclusion. [^1] KISPI: Ethical Committee of the Canton of Zurich, Switzerland (Decision numbers: 2017-00885, 2016-01019, 2017-00167). CHUV: Approved by the ethics committee of the Canton de Vaud (decision number CERVD 2021 00124). dHCP: Approved by the UK Health Research Authority (REC: 14/LO/1169), with written parental consent. For each subject, we used a super-resolution reconstructed T2-weighted MRI, resampled to an isotropic resolution of 0.5×0.5×0.5mm3 and padded or cropped to a fixed size of 256×256×256 voxels. For the KISPI, dHCP, and STA datasets, annotations were harmonized into a consistent 8-class labeling scheme: cerebrospinal fluid (CSF), grey matter (GM), white matter (WM), ventricles (VM), cerebellum (CBM), deep grey matter (SGM), brainstem (BSM), and CC. Details of the harmonization procedure are provided in the Supplementary Section S1.
@@
dHCP [28] Philips Achieva (3T) MB‐TSE TR/TE: 2265/250 ms 1.1×1.1×2.2 mm³ IRTK [26]; [29] $0.8^{3}$ 21–38 248/0 dHCP [30]
STA [31] Siemens Skyra / Trio (3T) Philips Achieva (1.5T) SS‐FSE TR/TE: 1400–2000/100–120 ms 0.9–1.1×0.9–1.1×2.0 mm³ SR1 [32] $1^{3}$ 21–38 18/0 FeTA [24]+ CC [27]
@@ 2.2 Pathology-Informed Domain Randomization for CCD
To ensure plausibility, an expert neuroradiologist reviewed the generated segmentations, both individual alterations and their combinations, to confirm broad anatomical constraints (for example, maintaining realistic ventricular dilation limits and proper continuity in simulations of pontocerebellar hypoplasia). However, the expert did not assess or correct finer‑grained details (such as subtle changes in adjacent structures like the cingulate gyrus), both because detailed segmentations of these regions were not available and because enforcing such precision would conflict with the goal of domain randomization.
@@ 4.0.1 Acknowledgements
Supplementary Material
S1 Annotation Harmonization
To ensure consistent training and evaluation across datasets with differing annotation protocols, we harmonized all labels to match the 7-class FeTA Challenge scheme [24], which includes cerebrospinal fluid (CSF), gray matter (GM), white matter (WM), ventricles (VM), deep gray matter (SGM), cerebellum (CBM), and brainstem (BSM). For the KISPI dataset, we used annotations from Fidon et al. [27], which refine the original FeTA labels and add a separate corpus callosum (CC) class. In the dHCP dataset [6], annotations were produced using the Draw-EM pipeline [30] and manually remapped to the FeTA classes following the approach in [17], with label mappings such as (Draw-EM → FeTA): 1→1 (CSF), 2→2 (GM), 3→3 (WM), 4→0 (background), 5→4 (LV), 6→5 (CBM), 7→6 (SGM), 8→7 (BSM), and 9→3 (merged with WM). Notably, this merges the hippocampi and amygdala (label 9) with WM, and the skull (label 4) with the background. Anatomical definitions also differ slightly between the two schemes—for example, Draw-EM includes only the lateral ventricles under the “ventricles” label, while FeTA includes all four ventricles. The corpus callosum label from Draw-EM was retained as an additional class. The STA dataset [31] underwent the same remapping ad dHCP and also included the CC class. As a result, all datasets were standardized to an 8-class annotation scheme, comprising the 7 FeTA tissue classes plus the corpus callosum.
S2 Pathology-informed Generation Details
S2.1 Metrics
Generalized DSC (gDSC) [36] applies class-specific weights to handle imbalance:
$gDSC=2\frac{\sum_{c}w_{c}\cdot|S_{g}^{c}~\cap~S_{t}^{c}|}{\sum_{c}w_{c}(|S_{g}^{c}|~+~|S_{t}^{c}|)}=\frac{2\sum_{c}w_{c}\cdot TP_{c}}{\sum_{c}w_{c}(2TP_{c}+FP_{c}+FN_{c})}$
where $w_{c}=1/(S_{g}^{c})^{2}$ is the inverse squared volume of class $c$.
Hausdorff Distance 95th Quantile (HD95) The Hausdorff Distance measures the largest boundary discrepancy between predicted and ground truth segmentations. To mitigate sensitivity to outliers, we use the 95th percentile version [37], defined as:
$HD_{95}(A,B)=\max(h_{95}(A,B),h_{95}(B,A))$
where
$h_{95}(A,B)=p_{95}\left(\min_{b\in B}\|a-b\|~\big|~a\in A\right)$
Distances are computed in voxels and reported in millimeters to allow comparisons across resolutions.
Volumetric Similarity (VS) VS compares the total volumes of the predicted and ground truth segmentations, regardless of spatial alignment:
$VS=1-\frac{||S_{t}|-|S_{g}||}{|S_{t}|+|S_{g}|}$
This is particularly useful for small structures like the corpus callosum, where spatial metrics may be unreliable due to segmentation noise.
S3 Results for KISPI-Path
Table S2 reports the Dice scores of all evaluated models on the KISPI‑Path dataset for each tissue class, including the corpus callosum, while Table S2 presents the corresponding HD95 scores. As shown in Table 3, the evaluated models exhibit only minor differences in their mean Dice and HD95 metrics. However, when comparing the average Euler Distance (ED) of the KISPI‑Path ground truth annotations (11.18±12.49) with the ED scores achieved by our models (Table 3), it becomes evident that our models produce segmentations with superior topological consistency. Among all models, SimPath50 achieves the lowest ED, highlighting the effectiveness of our augmentation pipeline in preserving anatomical topology.
Notably, when considering all metrics and tissue labels, SimPath50 consistently outperforms BSynth, which serves as a more relevant baseline since both rely on synthetic training data. These results demonstrate that our pathology‑informed augmentation strategy effectively reduces real‑synthetic domain gaps, leading to improved accuracy and robustness.
Table S1: Mean Dice scores ($\uparrow$) per structure across models.
Structure BReal BSynth SimPath50 SimPath100
CSF 0.68 ± 0.34 0.64 ± 0.33 0.64 ± 0.33 0.64 ± 0.34
GM 0.68 ± 0.22 0.61 ± 0.24 0.62 ± 0.22 0.61 ± 0.24
WM 0.85 ± 0.17 0.81 ± 0.18 0.83 ± 0.16 0.81 ± 0.18
VM 0.84 ± 0.16 0.78 ± 0.14 0.80 ± 0.12 0.80 ± 0.13
CBM 0.66 ± 0.39 0.78 ± 0.20 0.75 ± 0.22 0.77 ± 0.19
SGM 0.84 ± 0.14 0.79 ± 0.15 0.80 ± 0.12 0.80 ± 0.14
BS 0.66 ± 0.31 0.75 ± 0.15 0.73 ± 0.16 0.76 ± 0.12
$C$ 0.63 ± 0.15 0.56 ± 0.16 0.54 ± 0.14 0.44 ± 0.15
Table S2: Hausdorff distances ($m$, $\downarrow$) per structure across models.
Structure BReal BSynth SimPath50 SimPath100
CSF 6.78 ± 9.51 6.95 ± 9.23 7.09 ± 9.52 7.10 ± 9.32
GM 2.05 ± 2.22 1.70 ± 1.27 1.61 ± 1.06 1.65 ± 1.08
WM 3.23 ± 4.70 2.28 ± 1.46 2.03 ± 1.17 2.10 ± 1.28
VM 3.66 ± 4.28 2.29 ± 1.73 1.99 ± 1.09 2.46 ± 2.41
CBM 4.70 ± 6.66 2.70 ± 2.36 3.32 ± 3.51 4.27 ± 6.14
SGM 3.17 ± 4.89 2.70 ± 1.30 2.58 ± 0.82 2.74 ± 1.26
BS 6.04 ± 7.38 2.42 ± 1.10 2.87 ± 1.42 2.71 ± 1.94
$C$ 3.82 ± 5.76 3.04 ± 4.59 3.04 ± 4.26 5.06 ± 4.53